For the closed-loop unity feedback system given below Kp - 350, Kd = 20. Find the maximum overshoot Mp? 1 S(+2)
Answer #1

1 solno Gers). wa 350 un 18.7 radts controller gain= G(S) = (kp + Skd) Gels)= (350+208) GIS CC8) SCS+2) G(3) 1 → G0350+208)Ipage-2 maximum overshool Mp3 = evi-18.88772 SR e si-ył XlooY. -0.587T Mp Xlob 4 -2.27 x 100 m Mp -10-23%